The speed of light in a particular transparent medium is 2/8 the speed of light in a vacuum. What is the index of refraction of the material?

Group of answer choices

1.3

0.75

4

0.25
Answer #1

index of refraction of material is n= c/v

n = c/(2/8c) = 8/2 = 4

index of refraction for this material is n = 4